Upd a.com/af/servlet/quiz?ctx-adamghuloum-0092&quiz action-takeQuiz&quiz_probGuid-ONAPCOA8O upda 3. Gains from trade Consider two neighboring island countries, Arcadia and Dolorium. Each has 24,000 labor hours available that it carn use to produce jeans, wheat, or a combination of both. The following table shows the amount of labor hours required to produce one pair of jeans or one bushel of wheat. Jeans (Labor hours per pair) 48 20 Wheat (Labor hours per bushel) 12 10 Arcadia Dolorium has an absolute advantage in the production of jeans, and has an absolute advantage in the production of wheat. Initially, suppose Arcadia uses 18,000 hours of labor to produce jeans and 6,000 hours to produce wheat, while Dolorium uses 6,000 hours of labor to produce jeans and 18,000 hours to produce wheat. Consequently, Arcadia produces 375 pairs of jeans and 500 bushels of wheat, and Dolorium produces 300 pairs of jeans and 1,800 bushels of wheat. Assume there are no other countries willing to trade goods, so in the absence of trade between these two countries, each country consumes the amount of jeans and wheat it produces Arcadias opportunity cost of producing one pair of jeans is cost of producing one pair of jeans is advantage in the production of jeans, and of wheat, and Doloriums opportunity of wheat. Therefore, haş a comparative has a comparative advantage in the production of wheat. Suppose that Arcadia completely specializes in the production of the good in which it has a comparative advantage, producing only that good. It will produce Suppose also that Dolorium does not specialize and uses 18,000 hours of labor to produce jeans and 6,000 hours of labor to produce wheat. It will produce pairs of jeans and bushels of wheat.
Answer #1

Dolorium has an absolute advantage in the production of jeans, and Dolorium has an absolute advantage om the production of wheat. (lower labour hours required)

Arcadia's opportunity cost of producing one pair of jeans = 48/12 = 4 bushels of wheat, and Dolorium's opportunity cost of producing one pair of jeans = 20/10 = 2 bushels of wheat. Therefore, Dolorium has a comparative advantage in the production of jeans, and Arcadia has a comparative advantage in the production of wheat.

After specialization:

Arcadia will produce = 24000/12 = 2000 bushels of wheat

Dolorium will produce = 18000/20 = 900 pair of jeans and 6000/10 = 600 bushels of wheat.